okay here is the question...Ticket sales for the band concert totaled $1200. The adult tickets were $2 each. Student tickets were $1 each. There were twice as many adults tickets sold as student tickets. How many of each kind were sold?
i need to see how you got the answer

Answer by LilSkittleMd(119)

Let A=adult tickets
S=student tickets
First you need to set up your equations
2A+1S=1200
A=2S
To solve this use the substitution method and plug in what A is equal to
2(2S)+1S=1200
Distribute
4S+1S=1200
Combine like terms
5S=1200
Divide both sides by 5
S=240
Now plug in the S value into one of the original equations
A=2(240)
A=480
There were 240 student tickets sold and 480 adult tickets sold